OpenStudy (anonymous):

y+4=2/3(x+7) how do I put this in standard form

OpenStudy (amistre64):

get x and y on the same side

OpenStudy (anonymous):

i need an explanation

OpenStudy (amistre64):

that is an explanation ... not real sure how to explain it otherwise.

OpenStudy (amistre64):

how do you define standard form?

OpenStudy (anonymous):

how would I set it up?

OpenStudy (anonymous):

\[ax + by = c\]

OpenStudy (anonymous):

i have to put those numbers in slope intercept form, point slope form and standard form

OpenStudy (anonymous):

Point slope form is what's given. \(y - y_1 = m(x - x_1)\) Slope intercept form is \(y = mx + b\) Standard form is what I gave you above, \(ax + by = c\)

OpenStudy (anonymous):

i am stressed

OpenStudy (anonymous):

Why? I gave you the forms. It's just plugging in the numbers. Be careful and you'll get it.